Pat Is X Years old. KAtie Is 5 years younger than pat
Write katies age in terms of x:
The wsum of PAt's Age and katies age is 27
write and equation in x to represent this info solve your equation to find pats age

Answer by rothauserc(4718)

we are given Pat's age = x and Katie's age = x - 5
:
x + x-5 = 27
:
2x - 5 = 27
:
2x = 32
:
x = 16
:
*******************
Pat is 16 years old
*******************
: